# Bastani, Bastani, and McLaughlin

### v2

@@ 4 Simulations
We perform four simulation experiments exploring the efficacy of our approach. The first two show how predictive model noise deteriorates the performance of policies on the estimated frontier by varying the training size and complexity of the causal relationship between covariates and counterfactual outcomes. The later two compare a policy from the estimated frontier to policies generated by other common policy optimization approaches.
4.1 Setup
Our simulations assign each unit a hidden type which determines the mean and variance of their counterfactual outcomes. These types are connected to the unit’s covariates via a randomly generated classification problem. We generate training and test datasets by assigning treatments uniformly at randomly and generating counterfactual outcomes from the distribution of the unit’s type. We train a set of honest random forest regressors (one for each treatments) on the training dataset to provide counterfactual outcome estimates that can generate an estimate of the Pareto frontier. Then, we estimate the performance of these policies and evaluate $z$-scores for those estimates on the test dataset using an IPW estimator. Simulations were performed in Python.
Upon generation, each unit $n$ is assigned a type $\theta_{n}$ from one of 5 possible types which determines their counterfactual outcome distributions for each of 25 possible treatments. If unit $n$ is assigned treatment $t$, their outcome is drawn from a normal distribution whose mean and variance and determined by their type $Y_{n,t}\sim\mathcal{N}(\mu_{\theta_{n},t},\sigma^{2}_{\theta_{n},t})$. We generate the counterfactual mean and variance of the outcome in response to each treatment for all types $\theta$ and treatments $t$ as,
$\mu_{\theta,t}$ $\overset{iid}{\sim}\begin{cases}\mathcal{N}(1,10^{-4})&\text{w.p. }\ 0.1\\ \mathcal{N}(-1,10^{-4})&\text{w.p. }\ 0.1\\ \mathcal{N}(0,10^{-4})&\text{w.p. }\ 0.8\end{cases}$ $\sigma_{\theta,t}^{2}=1+3\mu_{\theta,t}^{2}.$
Essentially each treatment can be thought of as good, bad, or neutral for each type, with small perturbations added ensure all treatments do not impose the same conditional average treatment effect for any type. Good and bad treatments are rare (in expectation, there are 2.5 of each treatment per type) and the outcomes following them are noisier, making them more challenging to identify. They occur in equal proportions so that an RCT over all units generates a near 0 average treatment effect, while the optimal policy’s average treatment effect should be near 1. The goal of policy optimization in this case can be thought of as attempting to maximize the number of good treatments assigned while minimizing the number of bad treatments.
We generate units in covariate–type pairs using a random classification problem. To do this, we use scikit learn’s make_classification method, which first sets up a number of cluster centers for each type in a continuous covariate space and then generates units (in terms covariates) randomly around each clusters’ center. We generate units with 10 covariates which could be informative of type and 10 noninformative covariates with the number of clusters per type depending on the specific simulation.
To generate observational data for both training and testing, we generate treatments uniformly at randomly, assigning them a counterfactual outcome from the distribution of their given type–treatment. This allows us to generate an observational dataset $\mathcal{O}=\{(\pi_{1}^{o},X^{o}_{1},T^{o}_{1},Y^{o}_{1})\}$ of 2500 samples against which we run out of sample performance tests. It also allows us to create training datasets $\mathcal{T}$ of various sizes to explore how our approach depends on the quality of input estimates.
As input to estimate the Pareto frontier of policies, we train one honest random forest for each treatment observed in the training dataset $\mathcal{T}$. Honest random forests grow a set of unbiased regression trees which each predict the impact of a treatment according to a unit’s covariates, then average the trees’ predictions to estimate a counterfactual mean (Athey et al. 2019). We use random forests implemented by EconML’s RegressionForest method with 1,000 trees and honest set to True (ensuring unbiased estimates suitable for causal inference). We then extract counterfactual variance predictions according to the bag of bootstraps method detailed in Lu and Hardin (2021). When generating the entire frontier from IAPO, we present the performance of $\hat{\pi}^{*}(\xi)$ for 25 values of $\xi$ spread logarithmically in $\xi\in\left[\frac{1}{2500},40\right].$
4.2 Estimating the Pareto Frontier
We study how the estimated Pareto frontier varies with predictive model error as controlled by (1) the size of the training dataset, and (2) the complexity of the causal relationship between covariates and counterfactual outcomes. Figure 1 shows the estimated Pareto frontier as we vary the error of the prediction model. The left holds the complexity of the causal model fixed by generating 3 clusters of covariates per type and explores how additional training samples improve the quality of the estimated frontier. The right fixes the number of training samples at 50,000 and varies the complexity of the causal model by generating either 1 (Low), 3 (Medium), or 9 (High) covariate clusters per type. We ensure the test dataset is the same regardless of the causal model by first drawing the types in test dataset and generating their observed treatments and outcomes, then connecting them with a set of covariates once generated. In addition, we report the performance of the true frontier (i.e., where our predictive model is generated by an oracle) as well as its expected performance ex-ante generation of the test dataset. We mark the performance of the estimated expectation maximizing policy for each condition with an “x”.
Figure 1: The performance of the estimated Pareto frontier improves with the accuracy of the predictive model. We see this via increased performance of policies on the frontier as the size of the training dataset increases and decreased performance as the model complexity increases.
As expected, as the prediction model error decreases (either because the number of samples increases or the complexity of the causal model decreases), policies on the estimated Pareto Frontier perform better. More interestingly, prediction model error impacts performance along the frontier heterogeneously. For lower values of $\xi$, $\hat{\pi}^{*}(\xi)$ improves consistently, finding more significant policies with larger improvements near the observational policy than it could with worse information. On the other hand, improvements are more sporadic when $\xi$ is high and $\hat{\pi}^{*}(\xi)$ is close to the estimated expectation maximizing policy. We see that the combination of counterfactual estimation and IPW estimation introduces enough noise that the estimated expectation maximizing policy trained on 10,000 samples can outperform the one trained on 25,000 samples (and similarly for medium and high complexity models). This finding is further supported by the fact that the medium complexity estimate of the Pareto frontier in the right figure works off a similarly complex causal model to the left figure. For low $\xi$, the medium complexity estimate of the Pareto frontier generates policies with a $z$-score near 6, similar to the frontiers in the left figure with 25,000 and 100,000 samples (whereas medium complexity was trained on 50,000 samples). On the other hand, the estimated expectation maximizing policy of these policies in the left figure achieved a $z$-score near 4 with an improvement of 0.6, while the medium complexity case achieved a $z$-score under 2 with an improvement of 0.2.
In addition, prediction model error impacts the shape of the estimated Pareto frontier. The variance introduced to the IPW estimator by the realization of the test dataset’s treatments and outcomes can curve the estimated Pareto Frontier so that increasing $\xi$ can either increase the $z$-score or decrease the estimated improvement of the policy. We observe this most often for relatively high $\xi$, where evaluation is the most variable. Furthermore, higher error prediction models generate estimated Pareto frontiers that rapidly lose significance as the dual variable $\xi$ increases. The noisier the prediction model, the lower $\xi$ for which $\hat{\pi}^{*}(\xi)$ may begin to eliminate good treatments for some units, leading to this deterioration. As such, we recommend setting a more conservative value of $\xi$ than suggested in Section 3.2 when prediction model error is a large concern.
4.3 Comparison to Alternate Approaches
We compare the performance of a policy off the estimated Pareto frontier to other common policy optimization techniques. Provided sufficiently many training samples and a sufficiently simple causal model, these alternate methods can extract large significant policy gains. However, IAPO can still consistently extract statistically significant policy gains when others fail due to limited training samples or highly complex causal models.
Figure 2 compares the performance of a policy on the estimated Pareto frontier against policies generated by other common policy optimization methods in similar environments as the simulations of Section 4.2. We refer to the policy from the estimated Pareto frontier with $\xi=0.128$ (chosen following Section 3.2 with $\lambda=0.4$ and $Z_{\min}=2.5$) as IAPO (for Inference-Aware Policy Optimization). We chose $\lambda$ by consulting Figure 1 and observing that policies built off of predictive models with higher error struggled to reach this improvement. We compare the performance of IAPO to (1) the estimated expectation maximizing policy according to the honest random forests (the same predictive model we use with IAPO) (Athey et al. 2019), (2) a forest of policy trees designed for direct policy optimization (Athey and Wager 2021), and (3) a regularized weight-based approach to policy optimization called Norm-POEM (Swaminathan and Joachims 2015b). We implement policy forest using EconML’s DRPolicyForest method with 1,000 trees. We implement Norm-POEM using PyTorch and the L-BFGS optimizer with a variable step size set by Wolfe line search. Norm-POEM requires two hyper-parameters: $M$ (a variance at which to clip importance weights) and $\gamma$ (the regularization penalty). We run Norm-POEM for all combinations of $M\in\{1,\infty\}$ (with no weight inflation or no clipping) and 10 values of $\gamma$ spread logarithmically from 1 to 100 (the range in which the policy is responsive to $\gamma$ given these inputs), and report the most significant and best expected result for each predictive model. These policies often coincide, leading to an 8 sided star shape in the figure. We use smaller sample sizes in these simulations (10,000 training samples in the right figure) to focus on the setting where extracting a policy that is statistically significantly better than the observational policy is challenging.
Figure 2: Given sufficient training samples and a simple enough causal model, most policy optimization methods can extract significant policy improvements. However IAPO can do so consistently with fewer trainings samples and complex causal models.
IAPO can obtain significant policy improvements with less training samples and higher model complexities than other policy optimization techniques. At only 5000 training samples (or 10000 samples with a high causal model complexity), IAPO discovers a significant policy gain (at a 95% confidence level) while other techniques deliver policy improvements that fail a significance test. An exception to this is Norm-POEM, which may be achieving its policy gains due to either its focus on extrapolation or the advantage is receives being the best policy evaluation over many hyperparameter configurations. Furthermore, from Figure 1, we know significant policy improvements can be obtained using IAPO with only 2500 training samples and a smaller value of $\xi$. With more training samples or a sufficiently simple causal model, all techniques obtain statistically significant policy improvements, and IAPO begins to sacrifice expected policy improvement to deliver even more confidence for the specific tradeoff we chose. Additionally, IAPO delivers more consistent improvements than other policy optimization techniques; it’s policies deliver larger improvements as predictive model error decreases, whereas all other techniques have instances where more information about counterfactual outcomes lead to worse policy improvements, indicating high variability.
The right side of Figure 2 shows that the complexity of the causal model impacts the performance of Policy Forest and Norm-POEM substantially more than approaches based on honest random forests (such as IAPO). This finding may be because Policy Forest and Norm-POEM directly optimize policies over some policy class on the training dataset, which delivers more favorable regret bounds on policy optimization than plug-in regression approaches (Beygelzimer and Langford 2008), provided a near-optimal policy can be found within the policy class. Complex policy classes are difficult to compute, which is why Policy Forest and Norm-Poem use decision trees and exponential models, respectively. The optimal policies for complex causal models may not be easy to identify for these policy classes, leading to sub-optimal performance. By separating the problems of counterfactual prediction and policy optimization, plug-in approaches such as IAPO can identify complex causal relationships and suggest policies that exploit the confident predictions.
5 Technical Discussion
We discuss how prediction model errors impact the policies formed by Inference-Aware Policy Optimization (IAPO), and compare its plug-in approach to weight-based approaches.
5.1 The Impact of Prediction Errors
IAPO presents the optimal policy family for the conditional means and variances it is given, but the machine learning model’s predictions are subject to errors which hamper policy optimization (Qian and Murphy 2011). We can understand the impact of prediction errors on IAPO’s regularization by studying the path IAPO takes from the observational policy. A non-zero policy $\pi^{*}_{n,t}(\xi)$ changes in the direction
$\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}}\right),$
where for any $t\in T$, $\frac{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}{\pi_{n,t}^{o}}=\mathbb{E}\left[\left(Y_{n}^{o}\frac{\mathbf{1}\left[T_{n}^{o}=t\right]}{\pi^{o}_{n,T_{n}^{o}}}\right)^{2}\right]$ is the expectation of the squared importance weighted rewards under the policy where subject $n$ is given treatment $t$. This is a sufficient statistic and monotonically aligned with the amount of variance this deterministic policy would introduce on subject $n$ for any fixed mean. IAPO moves treatments slower if they contribute higher variance. Moreover, it only increases treatments if their expected outcome is above a weighted mean of the non-zero treatments’ expected outcomes that overweights low-variance contributions. Thus, IAPO still prefers policies in rank order of their expected outcome, but changes those with larger variance contributions more cautiously.
Despite these adjustments, IAPO’s policy prescriptions are still closely tied with the accuracy of its machine learning model, as seen in Figure 1. Obtaining policies with near-optimal improvements often requires IAPO to limit itself to positive policy weights on a subset of the treatments. This subset will only include the treatments with the highest expected outcomes and will thus be more vulnerable to estimation errors which reorder the perceived best treatments. Moreover, if any estimation error is accompanied by a low estimated variance, IAPO will quickly eliminate or promote that treatment, degenerating the quality of its policies, as seen in Figure 1 by the rapid significance deterioration of policies on frontiers estimated by particularly noisy prediction models. These problems are not specific to IAPO, but are more generally core issues faced when attempting to learn and evaluate near-optimal policies (Kitagawa and Tetenov 2018).
5.2 Plug-In Approaches vs. Weight-Based Approaches
We discuss tradeoffs between plug-in and weight-based approaches to policy learning. A shortcoming of plug-in approaches is that the counterfactual prediction model is typically trained using a standard objective such as mean-squared error (MSE) or cross-entropy loss, which may not be aligned with the downstream policy optimization objective. However, modern machine learning algorithms have become sufficiently expressive and powerful that they can form accurate predictions for highly complex inputs even for standard losses, thereby mitigating this issue. More broadly, plug-in approaches have the advantage that they automatically benefit from advances in supervised learning algorithms.
In contrast, weight-based approaches have two key shortcomings. First, the IPW-based objective estimate typically has high variance—many samples are effectively dropped due to receiving zero weight, thereby significantly reducing the effective size for policy estimation. There have been efforts to mitigate this problem by including regularization in the policy learning objective (Swaminathan and Joachims 2015a, b, Kallus 2018). A second shortcoming of weight-based approaches is the need to commit to a computationally tractable policy class. If the policy class is too simple, it will introduce biases such as eliminating heterogeneity for samples close in policy space (Stoye 2009), causing issues for complex causal models such as the one in Figure 2. Conversely, an overly complex policy class will be very difficult to optimize and sensitive to spurious patterns in the training data. In extreme cases, weight-based models can fit themselves to learn the mapping of covariates to treatments in the training data (Swaminathan and Joachims 2015b), producing a policy completely disconnected from the counterfactual outcomes distribution.

### v3

@@ 6 Simulations
We perform simulations exploring our method’s ability to extract policy improvements when naïve techniques fail. We apply our method to data from a megastudy encouraging flu vaccination using text message reminders (Milkman et al. 2022) and find a personalized policy that increases the vaccination rate by 4% while the naïve approach fails to extract a significantly improved policy (Section 6.1). Then, we simulate data based on the refugee relocation study of Bansak et al. (2018), showing our method can be effectively utilized in settings with additional policy constraints (Section 6.2). In Appendix 9 we provide additional simulations in a stylized environment in to explore how the accuracy of the prediction model given to our method impacts the performance of learned policies and compare our method to other policy optimization techniques.
6.1 Case Study: Vaccine Nudge Megastudy
Behavioral health interventions, such as text message reminders, offer a nearly costless mechanism for practitioners to improve health outcomes such as vaccination rates (Chapman et al. 2010, Milkman et al. 2011, Regan et al. 2017)—many people intend to get vaccinated, but fail to follow through (Sheeran 2002). Effective reminders can motivate individuals to act on their intentions, improving vaccination rates with minimal cost. We demonstrate that IAPO can be used to learn provably effective policies for targeting these kinds of nudges. We use data from Milkman et al. (2022), who performed a megastudy (i.e., a large-scale randomized control trial) testing the efficacy of 22 randomly chosen text messages and a control condition at encouraging 680,000 patients at a large US retail pharmacy to receive the flu vaccine during the 2019-2020 flu season.
The data includes the vaccination outcome of each patient (we use vaccinations before October 31, 2019) and their treatment assignment. In addition, it includes patient covariates related to their locations and health records, including (i) age, (ii) gender, (iii) zip code demographics, (iv) past vaccinations, and (v) hypertension. These covariates offer the opportunity to target the choice of text messages—different patients may respond to different motivations, offering the potential for large gains from a personalized treatment. We split the study subjects into training and test samples, train a set of honest random forests to predict each treatment’s counterfactual outcomes, use the forests to estimate the Pareto frontier of policies via IAPO, then estimate their performance on the test samples using IPW. After removing samples with missing covariates we maintain 474,667 samples for training and 203,429 samples for testing. In total 25.6% of the patients in our (combined training and test) sample received a vaccination by our cutoff.
Figure 2: The model believes policies on the estimated frontier can all achieve highly significant improvements; however a deterministic assignment of treatments following optimization of the model (marked by an ‘x’) does not achieve a confident improvement. IAPO still manages to extract significant policy improvements.
Figure 2 shows both the performance improvement the honest random forest model believes that policies on the estimated frontier can achieve (left) and the evaluation of those policies by IPW (right). It believes that naïvely choosing the “winning” treatment for each patient should yield statistically significant 20% increase in vaccination rate. However, when properly evaluated using IPW, this policy yields an improvement of 3.5%, and is not statistically significantly better. In contrast, IAPO learns a policy with a similar 3.5% increase in vaccination rate, which is statistically significantly better than the baseline policy at the 95% confidence level. Note that IAPO also suffers from the winner’s curse, since it is highly optimistic about performance along the frontier; however, it provides the regularization needed to obtain improvements that hold up under rigorous evaluation.
Table 1 describes how policies learned by IAPO achieve significant improvements when naïve optimization of our model fails. We focus on two policies from the frontier; 95% confident refers to a policy which achieves a similar improvement estimate to the naïve approach, but passes a significance test with 95% confidence while 99% confident refers to a policy which achieves half the improvement but passes a significance test with 99% confidence. All policies essentially eliminate the control treatment, but are highly personalized, with the most frequently used treatment given to less than 9% of patients. About half of the units are assigned a deterministic treatment by the 95% confident policy, while the 99% confident policy use stochastic treatments assignments for all units. This increases the overlap of both policies compared to the naïve policy, helping them achieve lower standard errors on their improvement estimates and allowing them to pass significance tests at higher confidence levels. The 95% confident policy shows that the naïve approach can be improved while only making half of units receive nondeterministic treatments while the 99% confident policy shows that much larger gains in significance require much greater overlap with the observational policy, assigning no treatments deterministically.
Policy Frequency (Min , Max) Active (Min, Mean, Max) Deterministic Overlap
Naïve (0.005 , 0.083) (1, 1, 1) 1.0 0.045
95% Confident (0.005 , 0.082) (1, 1.71, 7) 0.466 0.074
99% Confident (0.009 , 0.077) (2, 9.39, 19) 0.0 0.361
Table 1: Illustration of how IAPO policies differ from naïve ones. “Frequency” is the range of how frequently each treatment is assigned. “Active” is the number of non-zero treatment propensities for each unit on average. “Deterministic” is the fraction of units receiving a deterministic treatment assignment. “Overlap” is the similarity to the observational policy, reported as $1-TV(\pi,\pi^{o})$ where $TV$ is the total variation distance.
6.2 Case Study: Refugee Relocation Simulation
In the United States, refugee resettlement agencies are responsible for choosing where to settle incoming refugees. Their goal is to choose an assignment policy $\pi$ that places refugees in available locations in a welfare maximizing way, often measured by outcomes such as achieving employment within one year of settling. Assignment policies must not overload locations’ capacities (due to limited resources available to settle refugees at each location). We perform a simulation study to evaluate the efficacy of IAPO at policy learning when varying the best possible performance gains from personalization.
Specifically, we simulate a population of refugees matching the descriptive statistics of the population studied by Bansak et al. (2018). We generate 33,782 samples with characteristics of (i) location assigned, (ii) case status (free/fixed assignment), (iii) age, (iv) nationality, (v) education level, and (vi) arrival date. Each sample was assigned characteristics independently at random with probabilities matching the descriptive statistics given in Table S1 of Bansak et al. (2018). Age and nationality information were reported at a coarsened level, so we drew those characteristics uniformly at random from coarsened buckets. We refer to characteristics (iii)-(vi) of sample $n$ as $X_{n}$ and index different characteristics using subscript $j$. We use case status to determine whether samples are used for model training (24,327 samples) or testing (9,455 samples).
We simulate an employment outcome for each sample $n$ independently according to a bernoulli distribution whose mean $\mu_{n,t}$ depends on the sample’s characteristics $X_{n}$ and assigned location $t$. We standardize the characteristics $X_{n}$ to $X_{n}^{\dagger}$ by subtracting the average of each characteristic and dividing by that characteristic’s standard deviation across all samples. Then, we generate a fixed causal effect for each covariate $\alpha_{j}$ and each covariate-location interaction $\beta^{(t)}_{j}$; these interactions are what enable gains from personalization. We determine $\mu_{n,t}$ using a logistic function whose input is determined by a linear interaction of $\alpha$ and $\beta^{(t)}$ with $X_{n}^{\dagger}$:
$Y_{n,t}$ $\sim\mathrm{Ber}(\mu_{n,t})$ $\mu_{n,t}$ $=\frac{1}{1+e^{-X^{\dagger}_{n}\left(\alpha+\beta^{(t)}\right)+c_{tot}}}$ $\frac{\alpha_{j}}{c_{unit}}$ $\sim\mathrm{Unif}(-1,1)$ $\frac{\beta^{(t)}_{j}}{c_{intx}}$ $\sim\mathrm{Unif}(-1,1)$
The parameters $c_{unit}$ and $c_{intx}$ allow us to control the magnitude of the causal effects we introduce while $c_{tot}$ allows us to change the average employment outcome. Increasing $c_{unit}$ increases the heterogeneity in refugees’ baseline employment probabilities independent of location while increasing $c_{intx}$ increases the heterogeneity in the impact each location has on each refugee’s employment probability. Table 2 gives the parameters used to generate 3 datasets with varying magnitudes of covariate-location interaction effects. For each, we measure cross-unit heterogeneity as the standard deviation across samples employment probability (averaged over locations) and within-unit heterogeneity as the average standard deviation of samples employment probability across locations:
Cross-Unit Het. $=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\left(\overline{\mu}_{n}-\overline{\mu}\right)^{2}}$ Within-Unit Het. $=\frac{1}{N}\sum_{n=1}^{N}\sqrt{\frac{1}{K}\sum_{t=1}^{K}\left(\mu_{n,t}-\overline{\mu}_{n}\right)^{2}}$
$\overline{\mu}$ $=\frac{1}{N}\sum_{n=1}^{N}\overline{\mu}_{n}$ $\overline{\mu}_{n}$ $=\frac{1}{K}\sum_{t=1}^{K}\mu_{n,t}$
Our simulated refugee datasets maintain a consistent average outcome and cross-unit heterogeneity while varying the impact locations can have on a refugee’s employment.
Interaction $c_{tot}$ $c_{unit}$ $c_{intx}$ Employed Cross-Unit Het. Within-Unit Het.
None 0.86 0.25 0.0 0.316 0.136 0.0
Small 0.86 0.25 0.075 0.318 0.135 0.037
Medium 0.86 0.25 0.15 0.323 0.132 0.074
Table 2: Parameters used to generate eachof our three refugee simulation datasets with varying magnitudes of covariate-location interaction effects.
One challenge applying IAPO to this setting is the presence of resource constraints that limit the set of feasible refugee assignment policies. Each location $t$ can accept a maximum number of refugees $d_{t}$ given their available resources. We consider a policy feasible if, in expectation, it assigns fewer than $d_{t}$ refugees to location $t$, as represented by the constraint
$\sum_{n=1}^{N}\pi_{n,t}\leq d_{t}\ \forall\ t.$
Meeting the assignment constraints themselves can then be included as part of the randomization process or can be ignored if the constraints are approximate or somewhat flexible. For our simulations, we set each locations capacity $d_{t}$ as $5/4$ of the number of freely assigned refugees (our test samples) settled in location $t$. This choice allows proposed policies to slightly increase the number of refugees sent to any given location while requiring the majority of gains to come from identifying refugee-location synergy.
We integrate resource constraints into the IAPO pipeline as linear constraints to the convex optimization problems defining $\Pi_{\Lambda}$. Let $D=\left\{\pi\in(\Delta^{K})^{N}\mid\sum_{n=1}^{N}\pi_{n}\leq d\right\}$ be the set of feasible policies according to our capacities $d$. Then, our Pareto frontier is
$\Pi_{\text{Opt}}=\left\{\pi\in D\biggm|\not\exists\ \pi^{\prime}\ \in D\text{ s.t. }\tau(\pi^{\prime})\geq\tau(\pi)\wedge\frac{\tau(\pi^{\prime})}{s(\pi^{\prime})}\geq\frac{\tau(\pi)}{s(\pi)}\wedge(\tau(\pi^{\prime}),s(\pi^{\prime}))\neq(\tau(\pi),s(\pi))\right\},$
which can be relaxed to the family of convex optimization problems given by
$\Pi_{\Lambda}=\bigcup_{\lambda\in[0,\tau_{\text{max}}]}\operatorname*{\arg\min}_{\{\pi\in(\Delta^{K})^{N}\mid\tau(\pi)=\lambda,\sum_{n=1}^{N}\pi_{n}\leq d\}}s^{2}(\pi).$
Any policy constraint which can be represented as a valid convex programming constraint in terms of the policy’s propensities can be incorporated into IAPO in the same way as resource constraints.
Instead of characterizing the resource constrained $\Pi_{\Lambda}$, we approximate it using convex optimization software. Specifically, we implement the convex program defining $\Pi_{\Lambda}$ in CVXPY and solve it using the Splitting Conic Solver (SCS) while varying the value of $\lambda$ to solve for $\Pi_{\Lambda}$ point-wise. We use a convergence tolerance of $10^{-6}$ to achieve precise results; however, as our objective approaches zero for policies close to the observational policy, we only solve for policies with sufficiently high expected improvement $\lambda$ to ensure stability of the solver’s estimate.
Figure 3: Pareto frontier of policies computed using IAPO for each of our three refugee simulation datasets according to IPW estimation on a held out test set. The purple “x” marks the predicted performance of the expectation-maximizing policy, and the green “x” marks its actual performance.
We train a set of honest random forest regressions on each of our three refugee datasets to estimate the Pareto frontier of policies for each one. Figure 3 reports IPW performance estimates using held-out test sets, as well as the predicted and actual performance of the corresponding expectation-maximizing policies. For all three datasets, the winner’s curse causes the model to believe it can obtain massive employment gains regardless of the actual potential to improve employment. In contrast, when useful within-unit heterogeneity exists, IAPO obtains gains with much higher confidence and at least as much improvement as the naïve approach. In more detail, when policy improvement is impossible (“No Interaction Effect”), no policy along the frontier obtains significant policy improvement, which is expected. For the “Small Interaction Effect” dataset, the naïve approach fails to obtain statistically significant policy improvement, whereas IAPO obtains policy improvements up to the 99% confidence level while maintaining similar gains as the naïve approach. For the “Medium Interaction Effect” dataset, the naïve approach just barely obtains significant improvements at the 95% confidence level; in contrast, IAPO readily obtains similar improvements at a much higher 99% confidence level.